# Semi-classical imprints on quasinormal mode spectra

### v2

@@ I Introduction
Our procedure differs from previous work in this area in several important ways. First, semi-classical treatments often rely on simplifying approximations to the energy-momentum tensor (EMT), which has not been computed for generic fields in four dimensions. We work in a setting where no assumptions are made about the quantum state and all matter content is treated jointly, allowing us to provide general constraints at the expense of a precise characterization of the EMT and geometry. Other studies examining QNMs in quantum settings have focused on specific metrics which are tied to underlying assumptions about the correct quantization of gravity [19, 20, 21], limiting their applicability to those settings. And while Padé-like approximants have been employed previously in the study of quasinormal modes, namely in [22, 23], their approximation was designed to incorporate parameterized post-Newtonian (PPN) corrections, while we seek to isolate the contribution of near-horizon effects to the spectrum and present a more constrained form of the metric. To facilitate the interpolation we use a related, but distinct, M-fraction family of Padé approximants. Even in its preliminary form our scheme shows that, at least when dealing with restricted observational information, the QNMs of more general spherically-symmetric metrics have the potential to mimic some expected signatures of Kerr black holes. Finally, though stated in the context of semi-classical gravity, the method of analyzing QNMs we present here can be easily modified and applied to a wide variety of spacetimes in a purely classical setting.
@@ II Quasinormal modes
$f_{\ell m0}=f_{\ell m0}^{\mathrm{GR}}\left(1+\delta\hat{f}_{\ell m0}\right),$
@@ III Near-horizon metrics
What possibilities can quasinormal modes help to discriminate? In this work, the QNM analysis relies on several key assumptions of semi-classical gravity which we assume to be valid throughout the paper. First, regardless of the role that is played by quantum effects, classical geometric notions such as a metric and curvature, as well as various notions of horizons, are well-defined. Second, the motion of test particles and fields on such a background geometry is determined by the relevant classical equations of motion. Semi-classical general relativity is then completed by taking the renormalized expectation value of the EMT as the source for Einstein’s equations:
@@
Here, $G_{\mu\nu}=R_{\mu\nu}-\tfrac{1}{2}Rg_{\mu\nu}$ is the Einstein tensor, and the EMT in general depends on all quantum fields $\hat{\Phi}$ and their conjugate momenta $\hat{\Pi}$ as well on the classical metric $g_{\mu\nu}$. The evolution of the field state $|\psi{\rangle}$ is supposed to be governed by the Schrödinger equation with the Hamiltonian $\hat{H}[\hat{\Phi},\hat{\Pi};g]$. The EMT may contain curvature counterterms that arise from renormalisation, as well as the cosmological constant, or higher-order curvature terms of modified metric theories of gravity. Though the quantum contribution to the right-hand side is locally suppressed by a factor of $\hbar$ relative to the classical part of $T_{\mu\nu}$, there are a variety of settings in which these corrections can become large and affect both the geometry of a black hole and its formation dynamics.
One of our goals is to map the space of parameters which define the near-horizon geometry to the frequencies of QNMs. The self-consistent approach aims to obtain the information about this region and the EMT in it based on a minimal set of assumptions. Explicitly, we assume the validity of semi-classical physics, the formation of trapped regions of spacetime in finite time according to distant observers, and the regularity (in sense of having finite curvature scalars) of their boundary. On the other hand, no assumptions about the EMT, large scale structure of the spacetime, or the underlying quantum states are necessary. Under these conditions, the restriction of spherical symmetry allows for an exhaustive classification of the allowed near-horizon geometries [13]. We present the summary of the relevant facts and classification of admissible geometries in Appendix A.
@@
where $r$ is the areal radius and $f=1-2M(t,r)/r$ is a scalar function with $M$ being the Misner–Sharp–Hernandez (MSH) mass. If the spacetime contains a black hole, its outer boundary — the outer apparent horizon — is given by the largest (on a non-cosmological scale) real root of $f=0$, in all foliations that respect spherical symmetry. Admissible solutions are classified according to the scaling of the effective EMT components, such as $\tau_{t}{\vcentcolon=}e^{-2h}T_{tt}$ (Eqs. (59)–(61)), where $\tau_{t}\propto f^{k}$ on the approach to the horizon. Static limits of the dynamic solutions belong to the class $k=1$. Well-known solutions, such as the Reissner–Nordström, Bardeen, Hayward, and Simpson–Visser black holes belong to this class. To simplify the exposition we consider only solutions with $h\equiv 0$.
@@
$f(\tilde{x})\sim a_{1}\tilde{x}+a_{2}\tilde{x}^{2}+\mathcal{O}(\tilde{x}^{5/2})\ ,$ (14)
@@
$\frac{a_{1}}{r_{g}}=1-\frac{3}{4}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}\ ,\quad\frac{a_{2}}{r_{g}^{2}}=-1+\frac{15}{8}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}\ .$ (19)
It is easy to check which energy conditions [12] are satisfied for metrics of this type. The null energy condition at the horizon is violated for $a_{2}<-1/r_{g}^{2}$, and the weak energy condition is violated there if $E<0$, i. e. $a_{1}>1/r_{g}$.
@@ IV Boundary conditions and Padé approximants
which is (loosely) lower-bounded by neutron star compactness to be $\epsilon\gtrsim-0.4$ [23]. The terms of order $r^{-2}$ originate from deviations of the underlying theory from GR or from the presence of electric/magnetic charge. The first sub-leading term in the Bardeen model is of order $r^{-3}$, and in the Hayward model is of order $r^{-4}$.
@@ IV.1 Padé approximants
Clearly, any such rational function approximation has the potential to develop poles for certain choices of the parameters where the denominator vanishes, and the precise location of the poles will differ with the order of interpolation and number of terms included in the near and far expansions. These must be avoided if they occur outside of the horizon as it makes such expressions unsuitable for use in Eq. (4). We thus look for the minimal approximations which focus on the leading parameters $a_{1}$ and $a_{2}$ of the near-horizon expansion and are free of poles in a sufficiently large parameter space.
We also require that the approximation reduces to the Schwarzschild metric when the coefficients $a_{1}$ and $a_{2}$ take on their Schwarzschild values. This does not occur for certain otherwise pole-free Padeé approximations. Hence we consider four terms in the near-horizon expansion
$f(r)=\sum_{m=1}^{4}a_{m}\tilde{x}^{m}+\mathcal{O}(\tilde{x}^{5})\ ,$ (29)
which includes two subleading terms at orders $\tilde{x}^{3}$ and $\tilde{x}^{4}$ in the near-horizon expansion. This is the minimal number of additional terms required to construct a two-point Padé approximation which is well-behaved in the limit as $a_{1}\rightarrow 1/r_{g}$ and $a_{2}\rightarrow-1/r_{g}^{2}$ and does not develop poles in the region exterior to the horizon near this limit, at least for a finite domain in the parameter space $(a_{1},a_{2})$.
@@
$a_{3}=1/r_{g}^{3}\ ,\quad a_{4}=-1/r_{g}^{4}\ ,\quad M=r_{g}/2\quad(\epsilon=0)\ .$ (30)
There is no a priori reason to expect that if the parameters $(a_{1},a_{2})$ in Eq. (29) are different from their Schwarzschild counterparts by some $\delta_{a}$ (i.e. $|a_{1}r_{g}-1|\sim|a_{2}r_{g}^{2}+1|\sim\delta_{a}$) that the rest of the parameters $(a_{3},a_{4},\epsilon)$ can be taken equal to their Schwarzschild values. In fact, this is not true in many models. On the other hand, the results of Sec. VI show that $\delta_{a}\lesssim 10^{-2}$ already leads to differences between the QNM frequencies relative to those of the Schwarzschild metric of order $\Delta\omega\sim\mathcal{O}(1)$. However, taking the remaining parameters $(a_{3},a_{4},\epsilon)$ to differ from their Schwarzschild values also on the order of $\delta_{a}$, we find that the change in frequency is at most on the order of $\Delta\omega\times\delta_{a}$, justifying our approximation for the level of precision presented here.
@@
$A$ $=\left(3a_{1}^{2}r_{g}+2a_{1}a_{2}r_{g}^{2}-4a_{1}-3a_{2}r_{g}\right)$
$B$ $=\left(-17a_{1}^{2}r_{g}^{2}-12a_{1}a_{2}r_{g}^{3}+21a_{1}r_{g}+16a_{2}r_{g}^{2}\right)$
$C$ $=\left(37a_{1}^{2}r_{g}^{3}+28a_{1}a_{2}r_{g}^{4}-43a_{1}r_{g}^{2}-33a_{2}r_{g}^{3}+r_{g}\right)$
$D$ $=\left(-39a_{1}^{2}r_{g}^{4}-32a_{1}a_{2}r_{g}^{5}+44a_{1}r_{g}^{3}+34a_{2}r_{g}^{4}-3r_{g}^{2}\right)$
$E$ $=\left(20a_{1}^{2}r_{g}^{5}+18a_{1}a_{2}r_{g}^{6}-22a_{1}r_{g}^{4}-18a_{2}r_{g}^{5}+3r_{g}^{3}\right)$
$F$ $=(-4a_{1}^{2}r_{g}^{6}-4a_{1}a_{2}r_{g}^{7}+4a_{1}r_{g}^{5}+4a_{2}r_{g}^{6}-r_{g}^{4})$
@@
$\tilde{B}$ $=\left(-14a_{1}^{2}r_{g}^{2}-10a_{1}a_{2}r_{g}^{3}+17a_{1}r_{g}+13a_{2}r_{g}^{2}\right)$
$\tilde{C}$ $=\left(26a_{1}^{2}r_{g}^{3}+20a_{1}a_{2}r_{g}^{4}-30a_{1}r_{g}^{2}-23a_{2}r_{g}^{3}+r_{g}\right)$
$\tilde{D}$ $=\left(-24a_{1}^{2}r_{g}^{4}-20a_{1}a_{2}r_{g}^{5}+27a_{1}r_{g}^{3}+21a_{2}r_{g}^{4}-2r_{g}^{2}\right)$
$\tilde{E}$ $=\left(11a_{1}^{2}r_{g}^{5}+10a_{1}a_{2}r_{g}^{6}-12a_{1}r_{g}^{4}-10a_{2}r_{g}^{5}+2r_{g}^{3}\right)$
$\tilde{F}$ $=(-2a_{1}^{2}r_{g}^{6}-2a_{1}a_{2}r_{g}^{7}+2a_{1}r_{g}^{5}+2a_{2}r_{g}^{6})$
The algorithms for obtaining this expansion are described in [35]. Appendix B illustrates the relationship between coefficients in the continued fraction representation and the rationalized function for $n=3$. The expansion of $f(r)$ as given by (31) about $r=r_{g}$ and $r=\infty$ can easily be verified to coincide with (29) and (20) where the parameters are given by Eq. (30). Lower-order ($n<4$) approximations either do not agree in their expansions up to the given order or develop poles on the interval $r\in(r_{g},\infty)$ making them ill-suited for numerical integration. The $n=5$ approximation is chosen for its convenient domain of validity in the parameter space of $(a_{1},a_{2})$. In Figure1 and Figure2 we show, respectively, the interpolating function along with the effective potential as they deviate from the Schwarzschild limit for some values of $a_{1}$ and $a_{2}$. Fig. Figure3 shows the domain in the $(a_{1},a_{2})$ plane of parameters for which both the $n=4$ and $n=5$ Padé approximations are pole-free.
Figure 1: Two-point Padé interpolation (31) of the near-horizon metric (29) and the asymptotic Schwarzschild geometry (20) with $a_{1}=1.145$ and $a_{2}=-1.076$ and $r_{g}=1$. The horizon is located at the red dashed line.
Figure 2: Behaviour of the effective potential as the coefficients $a_{1}$ and $a_{2}$ in (14) deviate from their Schwarzschild values. The gravitational radius is $r_{g}=1$.
Figure 3: Parameter space for the two-point Padé approximation $f(r)$ (yellow) and $g(r)$ (orange). Poles on the interval $r\in(r_{g},\infty)$ develop for coefficients lying in the shaded region. The gravitational radius is $r_{g}=1$. The Bardeen metric has $a_{2}\approx\tfrac{3}{2}-\tfrac{5}{2}a_{1}$ with $a_{1}\leqslant 1$ while the static Hayward metric has $a_{2}\approx\tfrac{5}{2}-\tfrac{7}{2}a_{1}$ with $a_{1}\leqslant 1$. The weak energy condition may be satisfied for $a_{1}\leqslant 1$. The null energy condition is violated for $a_{2}<-1$.
@@
$A$ $=(a_{1}^{2}r_{g}-3a_{1}-2a_{2}r_{g})$
$B$ $=(1+13a_{1}r_{g}-5a_{1}^{2}r_{g}^{2}+9a_{2}r_{g}^{2})$
$C$ $=(-3r_{g}-20a_{1}r_{g}^{2}+9a_{1}^{2}r_{g}^{3}-14a_{2}r_{g}^{3})$
$D$ $=(3r_{g}^{2}+14a_{1}r_{g}^{3}-7a_{1}^{2}r_{g}^{4}+9a_{2}r_{g}^{4})$
$E$ $=(-r_{g}^{3}-4a_{1}r_{g}^{4}+2a_{1}^{2}r_{g}^{5}-2a_{2}r_{g}^{5})$
$\tilde{B}$ $=(1-4a_{1}^{2}r_{g}^{2}+10a_{1}r_{g}+7a_{2}r_{g}^{2})$
$\tilde{C}$ $=(6a_{1}^{2}r_{g}^{3}-13a_{1}r_{g}^{2}-9a_{2}r_{g}^{3}-2r_{g})$
$\tilde{D}$ $=(-4a_{1}^{2}r_{g}^{4}+8a_{1}r_{g}^{3}+5a_{2}r_{g}^{4}+2r_{g}^{2})$
$\tilde{E}$ $=(a_{1}^{2}r_{g}^{5}-2a_{1}r_{g}^{4}-a_{2}r_{g}^{5})\ ,$
and compare the relative size of the difference between the two functions for a given choice of $a_{1}$ and $a_{2}$ with the deviation in frequency induced by those values. For the interpolating function $g(r)$, poles appear for different values of the coefficients compared to $f(r)$. Computation using both interpolations for the parameters $(a_{1},a_{2})$ belonging to the overlapping pole-free domains indicate that the results differ only by order $\delta_{a}^{2}$, and thus the $\mathcal{O}(1)$ deviations observed when $(a_{1},a_{2})$ vary are not an artifact of the choice of interpolation order.

### v3

@@ I Introduction
Our procedure differs from previous work in this area in several important ways. First, semi-classical treatments often rely on simplifying approximations to the energy-momentum tensor (EMT), which has not been computed for generic fields in four dimensions. We work in a setting where no assumptions are made about the quantum state and all matter content is treated jointly, allowing us to provide general constraints at the expense of a precise characterization of the EMT and geometry. Other studies examining QNMs in quantum settings have focused on specific metrics which are tied to underlying assumptions about the correct quantization of gravity [19, 20, 21], limiting their applicability to those settings. And while Padé-like approximants have been employed previously in the study of quasinormal modes, namely in [22, 23], their approximation was designed to incorporate parameterized post-Newtonian (PPN) corrections, while we seek to isolate the contribution of near-horizon effects to the spectrum and present a more constrained form of the metric. To facilitate the interpolation we use a related, but distinct, M-fraction family of Padé approximants. Even in its preliminary form our scheme shows that, at least when dealing with restricted observational information, the QNMs of more general spherically-symmetric metrics have the potential to mimic the expected signatures of Kerr black holes. Finally, though stated in the context of semi-classical gravity, the method of analyzing QNMs presented here can be easily modified and applied to a wide variety of spacetimes in a purely classical setting.
@@ II Quasinormal modes
$f_{\ell m0}=f_{\ell m0}^{\mathrm{GR}}(1+\delta\hat{f}_{\ell m0}),$
@@ III Near-horizon metrics
In this work, the QNM analysis relies on several key assumptions of semi-classical gravity which we assume to be valid throughout the paper. First, regardless of the role that is played by quantum effects, classical geometric notions such as a metric and curvature, as well as various notions of horizons, are well-defined. Second, the motion of test particles and fields on such a background geometry is determined by the relevant classical equations of motion.
Semi-classical general relativity is then completed by taking the renormalized expectation value of the EMT as the source for Einstein’s equations:
@@
Here $G_{\mu\nu}{\vcentcolon=}R_{\mu\nu}-\tfrac{1}{2}Rg_{\mu\nu}$ is the Einstein tensor, and the EMT in general depends on all quantum fields $\hat{\Phi}$ and their conjugate momenta $\hat{\Pi}$ as well on the classical metric $g_{\mu\nu}$. The evolution of the field state $|\psi{\rangle}$ is supposed to be governed by the Schrödinger equation with the Hamiltonian $\hat{H}[\hat{\Phi},\hat{\Pi};g]$. The EMT may contain curvature counterterms that arise from renormalisation, as well as the cosmological constant, or higher-order curvature terms of modified metric theories of gravity. Though the quantum contribution to the right-hand side is locally suppressed by a factor of $\hbar$ relative to the classical part of $T_{\mu\nu}$, there are a variety of settings in which these corrections can become large and affect both the geometry of a black hole and its formation dynamics.
One of our goals is to map the space of parameters which define the near-horizon geometry to the frequencies of QNMs, with the goal of being maximally model-independent and sensitive to the largest possible domain for each of the defining parameters. The methods we describe also represent a first step towards extracting the observational signatures of the metrics that are obtained within the self-consistent approach to black hole horizons [13]. This approach aims to obtain the information about the near-horizon region and the EMT in it based on a minimal set of assumptions. Explicitly, we assume the validity of semi-classical physics as defined above, and supplement it by the requirements of horizon observability and weak regularity. The former is expressed as a requirement of formation of trapped regions of spacetime in finite time according to distant observers. The weak regularity demands only that curvature scalars that are built by using algebraic operations on the Riemann tensor components are finite on their boundary. On the other hand, no assumptions about the EMT, large scale structure of the spacetime, or the underlying quantum states are necessary. Under these conditions, the restriction of spherical symmetry allows for an exhaustive classification of the allowed near-horizon geometries [13]. We present the summary of the relevant facts and classification of admissible geometries in Appendix A.
@@
where $r$ is the areal radius and $f=1-2M(t,r)/r$ is a scalar function with $M$ being the Misner–Sharp–Hernandez (MSH) mass. If the spacetime contains a black hole, its outer boundary — the outer apparent horizon — is given by the largest (on a non-cosmological scale) real root of $f=0$, in all foliations that respect spherical symmetry. Admissible solutions are classified according to the scaling of the effective EMT components, such as $\tau_{t}{\vcentcolon=}e^{-2h}T_{tt}$ (Eqs. (59)–(61)), where $\tau_{t}\propto f^{k}$ on the approach to the horizon. Static limits of the dynamic solutions belong to the class $k=1$. Well-known solutions, such as the Reissner–Nordström, Bardeen, Hayward, and Simpson–Visser black holes belong to this class. Many popular models (including these) assume $h\equiv 0$, and to simplify the exposition in the following we only consider such solutions.
@@
$f(\tilde{x})\sim\alpha_{1}\frac{\tilde{x}}{r_{g}}+\alpha_{2}\frac{\tilde{x}^{2}}{r_{g}^{2}}+\mathcal{O}(\tilde{x}^{5/2})\ ,$ (14)
@@
$\alpha_{1}=1-\frac{3}{4}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}\ ,\quad\alpha_{2}=-1+\frac{15}{8}\lambda^{2}+\mathcal{O}\big{(}\lambda^{4}\big{)}\ .$ (19)
It is easy to check which energy conditions [12] are satisfied for metrics of this type. The null energy condition at the horizon is violated for $\alpha_{2}<-1$, and the weak energy condition is violated there if $E<0$, i. e. $\alpha_{1}>1$.
@@ IV Boundary conditions and Padé approximants
which is (loosely) lower-bounded by neutron star compactness to be $\epsilon\gtrsim-0.4$ [23]. Both Schwarzschild and Bardeen black holes have $\epsilon=0$.
The terms of order $r^{-2}$ originate from deviations of the underlying theory from GR or from the presence of electric/magnetic charge. The first sub-leading term in the Bardeen model is of order $r^{-3}$, and in the Hayward model is of order $r^{-4}$.
@@ IV.1 Padé approximants
The well-known Rezolla-Zhidenko parameterization [22] is based on a different family of Padé approximants known as C-fractions. Their scheme is both philosophically and technically distinct from our own and we summarize these differences in Appendix B, while also presenting comparisons for some simple families of metrics at the end of this Section.
Clearly, any such rational function approximation has the potential to develop poles for certain choices of the parameters where the denominator vanishes, and the precise location of the poles will differ with the order of interpolation and number of terms included in the near and far expansions. These must be avoided if they occur outside of the horizon as it makes such expressions unsuitable for use in Eq. (4). We thus look for the minimal approximations which focus on the leading parameters $\alpha_{1}$ and $\alpha_{2}$ of the near-horizon expansion and are free of poles in a sufficiently large parameter space.
We also require that the approximation reduces to the Schwarzschild metric when the coefficients $\alpha_{1}$ and $\alpha_{2}$ take on their Schwarzschild values. This does not occur for certain otherwise pole-free Padeé approximations. Hence we consider four terms in the near-horizon expansion
$f(r)=\sum_{m=1}^{4}\alpha_{m}\left(\frac{\tilde{x}}{r_{g}}\right)^{m}+\mathcal{O}(\tilde{x}^{5})\ ,$ (29)
which includes two subleading terms at orders $\tilde{x}^{3}$ and $\tilde{x}^{4}$ in the near-horizon expansion. This is the minimal number of additional terms required to construct a two-point Padé approximation which is well-behaved in the limit as $\alpha_{1}\rightarrow 1$ and $\alpha_{2}\rightarrow-1$ and does not develop poles in the region exterior to the horizon near this limit, at least for a finite domain in the parameter space $(\alpha_{1},\alpha_{2})$.
@@
$\alpha_{3}=1\ ,\quad\alpha_{4}=-1\ ,\quad M=r_{g}/2\quad(\epsilon=0)\ .$ (30)
There is no a priori reason to expect that if the parameters $(\alpha_{1},\alpha_{2})$ in Eq. (29) are different from their Schwarzschild counterparts by some $\delta_{a}$ (i.e. $|\alpha_{1}-1|\sim|\alpha_{2}+1|\sim\delta_{a}$) that the rest of the parameters $(\alpha_{3},\alpha_{4},\epsilon)$ can be taken equal to their Schwarzschild values. In fact, this is not true in many models. On the other hand, the results of Sec. VI show that $\delta_{a}\lesssim 10^{-2}$ already leads to differences between the QNM frequencies relative to those of the Schwarzschild metric of order $\Delta\omega\sim\mathcal{O}(1)$. However, taking the remaining parameters $(\alpha_{3},\alpha_{4},\epsilon)$ to differ from their Schwarzschild values also on the order of $\delta_{a}$, we find that the change in frequency is at most on the order of $\Delta\omega\times\delta_{a}$, justifying our approximation for the level of precision presented here.
@@
$A$ $=\left(3\;\!\alpha_{1}^{2}+2\;\!\alpha_{1}\>\!\alpha_{2}-4\;\!\alpha_{1}-3\;\!\alpha_{2}\right)$
$B$ $=\left(21\;\!\alpha_{1}r_{g}-17\;\!\alpha_{1}^{2}r_{g}-12\>\!\alpha_{2}\;\!\alpha_{1}r_{g}+16\>\!\alpha_{2}r_{g}\right)$
$C$ $=\left(37\;\!\alpha_{1}^{2}r_{g}^{2}-43\;\!\alpha_{1}r_{g}^{2}+28\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{2}-33\;\!\alpha_{2}r_{g}^{2}+r_{g}^{2}\right)$
$D$ $=\left(44\;\!\alpha_{1}r_{g}^{3}-39\;\!\alpha_{1}^{2}r_{g}^{3}-32\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{3}+34\;\!\alpha_{2}r_{g}^{3}-3r_{g}^{3}\right)$
$E$ $=\left(20\;\!\alpha_{1}^{2}r_{g}^{4}-22\;\!\alpha_{1}r_{g}^{4}+18\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{4}-18\>\!\alpha_{2}r_{g}^{4}+3r_{g}^{4}\right)$
$F$ $=(4\;\!\alpha_{1}r_{g}^{5}-4\;\!\alpha_{1}^{2}r_{g}^{5}-4\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{5}+4\>\!\alpha_{2}r_{g}^{5}-r_{g}^{5})$
@@
$\tilde{B}$ $=\left(17\;\!\alpha_{1}r_{g}-14\;\!\alpha_{1}^{2}r_{g}-10\;\!\alpha_{2}\;\!\alpha_{1}r_{g}+13\>\!\alpha_{2}r_{g}\right)$
$\tilde{C}$ $=\left(26\;\!\alpha_{1}^{2}r_{g}^{2}-30\;\!\alpha_{1}r_{g}^{2}+20\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{2}-23\;\!\alpha_{2}r_{g}^{2}+r_{g}^{2}\right)$
$\tilde{D}$ $=\left(27\;\!\alpha_{1}r_{g}^{3}-24\;\!\alpha_{1}^{2}r_{g}^{3}-20\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{3}+21\;\!\alpha_{2}r_{g}^{3}-2r_{g}^{3}\right)$
$\tilde{E}$ $=\left(11\;\!\alpha_{1}^{2}r_{g}^{4}-12\;\!\alpha_{1}r_{g}^{4}+10\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{4}-10\;\!\alpha_{2}r_{g}^{4}+2r_{g}^{4}\right)$
$\tilde{F}$ $=(2\;\!\alpha_{1}r_{g}^{5}-2\;\!\alpha_{1}^{2}r_{g}^{5}-2\;\!\alpha_{1}\>\!\alpha_{2}r_{g}^{5}+2\>\!\alpha_{2}r_{g}^{5})$
The algorithms for obtaining this expansion are described in [35]. Appendix B illustrates the relationship between coefficients in the continued fraction representation and the rationalized function for $n=3$. The expansion of $f(r)$ as given by (31) about $r=r_{g}$ and $r=\infty$ can easily be verified to coincide with (29) and (20) where the parameters are given by Eq. (30). Lower-order ($n<4$) approximations either do not agree in their expansions up to the given order or develop poles on the interval $r\in(r_{g},\infty)$ making them ill-suited for numerical integration. The $n=5$ approximation is chosen for its convenient domain of validity in the parameter space of $(\alpha_{1},\alpha_{2})$. In Figure1 and Figure2 we show, respectively, the interpolating function along with the effective potential as they deviate from the Schwarzschild limit for some values of $\alpha_{1}$ and $\alpha_{2}$.  Figure3 shows the domain in the $(\alpha_{1},\alpha_{2})$ plane of parameters for which both the $n=4$ and $n=5$ Padé approximations are pole-free.
Figure 1: Two-point Padé interpolation (31) of the near-horizon metric (29) and the asymptotic Schwarzschild geometry (20) with $\alpha_{1}=1.145$ and $\alpha_{2}=-1.076$. The horizon is located at the red dashed line.
Figure 2: Behaviour of the effective potential as the coefficients $\alpha_{1}$ and $\alpha_{2}$ in (14) deviate from their Schwarzschild values.
Figure 3: The parameter space $(\alpha_{1},\alpha_{2})$ covered by the two-point Padé approximation $f_{5}(r)$ (yellow) and $f_{4}(r)$ (orange). It is assumed that $\alpha_{3}=1,\alpha_{4}=-1$ and $\epsilon=0$. Poles on the interval $r\in(r_{g},\infty)$ develop for coefficients lying in the shaded region. The Bardeen metric has $\alpha_{2}\approx\tfrac{3}{2}-\tfrac{5}{2}\alpha_{1}$ with $\alpha_{1}\leqslant 1$ while the static Hayward metric has $\alpha_{2}\approx\tfrac{5}{2}-\tfrac{7}{2}\alpha_{1}$ with $\alpha_{1}\leqslant 1$. The weak energy condition may be satisfied for $\alpha_{1}\leqslant 1$. The null energy condition is violated for $\alpha_{2}<-1$.
Figure 4: Parameter space $(\alpha_{1},\alpha_{2})$ as covered by the two-point Padé approximation using the Rezzolla–Zhidenko scheme (Appendix B with $n=2$ (yellow) and $n=3$ (orange). It is assumed that $\alpha_{3}=1,\alpha_{4}=-1$ and $\epsilon=0$. Poles on the interval $r\in(r_{g},\infty)$ develop for coefficients lying in the shaded region. The weak energy condition may be satisfied for $\alpha_{1}\leqslant 1$. The null energy condition is violated for $\alpha_{2}<-1$.
@@
$A$ $=(\alpha_{1}^{2}-3\>\!\alpha_{1}-2\>\!\alpha_{2})$
$B$ $=(r_{g}-5\>\!\alpha_{1}^{2}r_{g}+13\>\!\alpha_{1}r_{g}+9\>\!\alpha_{2}r_{g})$
$C$ $=(9\>\!\alpha_{1}^{2}r_{g}^{2}-20\>\!\alpha_{1}r_{g}^{2}-14\>\!\alpha_{2}r_{g}^{2}-3\>\!r_{g}^{2})$
$D$ $=(3\>\!r_{g}^{3}-7\>\!\alpha_{1}^{2}r_{g}^{3}+14\>\!\alpha_{1}r_{g}^{3}+9\>\!\alpha_{2}r_{g}^{3})$
$E$ $=(2\>\!\alpha_{1}^{2}r_{g}^{4}-4\>\!\alpha_{1}r_{g}^{4}-r_{g}^{4}-2\>\!\alpha_{2}r_{g}^{4})$
$\tilde{B}$ $=(r_{g}-4\>\!\alpha_{1}^{2}r_{g}+10\>\!\alpha_{1}r_{g}+7\>\!\alpha_{2}r_{g})$
$\tilde{C}$ $=(6\>\!\alpha_{1}^{2}r_{g}^{2}-13\>\!\alpha_{1}r_{g}^{2}-9\>\!\alpha_{2}r_{g}^{2}-2\>\!r_{g}^{2})$
$\tilde{D}$ $=(2\>\!r_{g}^{3}-4\>\!\alpha_{1}^{2}r_{g}^{3}+8\>\!\alpha_{1}r_{g}^{3}+5\>\!\alpha_{2}r_{g}^{3})$
$\tilde{E}$ $=(\alpha_{1}^{2}r_{g}^{4}-2\>\!\alpha_{1}r_{g}^{4}+\alpha_{2}r_{g}^{4})\ ,$
and compare the relative size of the difference between the two functions for a given choice of $\alpha_{1}$ and $\alpha_{2}$ with the deviation in frequency induced by those values. For the interpolating function $f_{4}(r)$, poles appear for different values of the coefficients compared to $f_{5}(r)$. Computation using both interpolations for the parameters $(\alpha_{1},\alpha_{2})$ belonging to the overlapping pole-free domains indicate that the results differ only by order $\delta_{a}^{2}$, and thus the $\mathcal{O}(1)$ deviations observed when $(\alpha_{1},\alpha_{2})$ vary are not an artifact of the choice of interpolation order.